OpenStudy (anonymous):

Select the equations that are parallel and perpendicular to y = 1/4x + 7 and that pass through the point (-8, 9).

OpenStudy (anonymous):

need options for this one

OpenStudy (anonymous):

parallel: y = one over four x + 11 perpendicular: y = -4x - 23 parallel: y = 4x - 7 perpendicular: y = negative one over four x - 7 parallel: y = one over four x - 11 perpendicular: y = -4x + 23 parallel: y = -4x + 11 perpendicular: y = one over four x - 4

OpenStudy (anonymous):

A